Circular No. 8822

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Mailstop 18, Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A.

IAUSUBS@CFA.HARVARD.EDU or FAX 617-495-7231 (subscriptions) CBAT@CFA.HARVARD.EDU (science)

        URL http://cfa-www.harvard.edu/iau/cbat.html ISSN 0081-0304

Phone 617-495-7440/7244/7444 (for emergency use only)

SUPERNOVAE 2007ag–2007as Thirteen apparent supernovae have been discovered on unfiltered CCD

frames and reported to the Central Bureau: 2007ag, 2007ao, and 2007ap by T. Puckett (cf. IAUC 8814; 2007ag and 2007ao with R. Gagliano, and 2007ap with A. Kroes); 2007an by M. Migliardi (via A. Dimai; cf. IAUC 8667); 2007ar by G. Duszanowicz (cf. IAUC 8789); 2007as by P. Luckas, O. Trondal, and M. Schwartz (cf. IAUC 8789); and the rest by W. Li (cf. IAUC 8813; 2007aj–2007al with D. R. Madison, 2007am with N. Joubert, and 2007aq with D. Winslow). Discovery observations:

SN  2007 UT  α_{2000}  δ_{2000}  Mag.  Offset
2007ag Mar. 7.29 10^{h}01^{m}35.^{s}99 +21^{o}36′42′′.0 18.0 4′′.0 E, 15′′.5 N
2007ah Mar. 6.13 4 02 18.50 +25 49 06.8 17.8 0′′.4 W, 4′′.9 N
2007ai Mar. 6.53 16 12 53.74 −21 37 48.7 18.0 5′′.1 E, 24′′.5 S
2007aj Mar. 3.43 12 47 54.45 +54 00 37.2 16.4 130′′ W, 47′′ S
2007ak Mar. 10.16 5 20 40.75 + 8 48 16.0 18.2 15′′.1 S
2007al Mar. 10.28 9 59 18.48 −19 28 25.8 16.6 3′′.9 W, 2′′.0 S
2007am Mar. 11.36 10 46 33.62 +13 45 09.3 17.8 20′′.1 W, 7′′.3 N
2007an Mar. 10.98 11 58 43.72 +27 27 01.2 17.5 24′′ W, 7′′ S
2007ao Mar. 13.38 14 16 49.74 +10 48 24.1 17.7 47′′ W, 2′′.4 S
2007ap Mar. 13.45 15 56 23.06 +16 30 57.9 15.5 8′′.0 W, 26′′.3 S
2007aq Mar. 13.21 8 48 22.03 +18 19 32.6 18.0 38′′ W, 20′′ S
2007ar Mar. 12.92 13 21 01.93 +58 33 02.4 17.5 7′′.4 W, 7′′.6 N
2007as Mar. 13.61 9 27 36.01 −80 10 39.2 15.8 12′′.1 W, 8′′.8 N

Additional information is available on the following Electronic Telegrams: 2007ag in UGC 5392, CBET 868 and 874 (type-Ib, discovered near maximum light); 2007ah in UGC 2931, CBETs 869 and 878 (type-II, discovered just past explosion); 2007ai in MCG –04-38-4, CBETs 870 and 880 (type-Ia, discovered just before maximum); 2007aj, CBETs 871 and 873 (type-Ia, discovered near maximum); 2007ak in UGC 3293, CBETs 875 and 881 (type-IIn); 2007al, CBETs 875 and 878 (type-Ia, discovered near maximum); 2007am in NGC 3367, CBETs 877 and 884 (type-II, discovered ∼ 10 days past explosion); 2007an in NGC 4017, CBETs 882 and 896 (type-II); 2007ao in NGC 5532, CBETs 883 and 892 (subluminous type-Ia, discovered ∼ 10 days past peak); 2007ap in MCG +03-41-3, CBET 883; 2007aq in IC 2409, CBETs 885 and 889 (type-II, discovered ∼ 3.5 weeks past explosion); 2007ar in MCG +10-19-62, CBETs 886 and 889 (type-Ia, discovered just past maximum); 2007as in ESO 18-G18, CBET 888.

2007 March 19  \begin{matrix}c \\ ◯ \end{matrix} Copyright 2007 CBAT  Daniel W. E. Green